Find the slope of the line(2,5) and (4,4).

Answer by Maths68(1474)

Find the slope of the line passing through the points (2,5) and (4,4).

Put the values of above points in the equation of straight line point-slope form
m=(y2-y1)/(x2-x1)
m=(4-5)/(4-2)
m=-1/2
Slope of the line = m = -1/2